OpenStudy (jamesreed30):

Express the required measurement in the exact form what is the degree measure of an ark for pie feet long in a circle of radius 10 feet

OpenStudy (jamesreed30):

4 Pi feet long*

OpenStudy (osprey):

done a sketch and encl as pptx it may help ....

OpenStudy (dumbcow):

arc length = angle*radius \[\pi = 10 \theta\] \[\theta = \frac{\pi}{10}\] Then you simply have to convert that to degrees (pi = 180)